# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ Abstract
We analyze the generalization gap (gap between the training and test errors) when training a potentially over-parametrized model using a Markovian stochastic training algorithm, initialized from some distribution $\theta_{0}\sim p_{0}$. We focus on Langevin dynamics with a positive temperature $\beta^{-1}$, i.e. gradient descent on a training loss $L$ with infinitesimal step size, perturbed with $\beta^{-1}$-variances Gaussian noise, and lightly regularized or bounded. There, we bound the generalization gap, at any time during training, by $\sqrt{(\beta\mathbb{E}L(\theta_{0})+\log(1/\delta))/N}$ with probability $1-\delta$ over the dataset, where $N$ is the sample size, and $\mathbb{E}L(\theta_{0})=O(1)$ with standard initialization scaling. In contrast to previous guarantees, we have no dependence on either training time or reliance on mixing, nor a dependence on dimensionality, gradient norms, or any other properties of the loss or model. This guarantee follows from a general analysis of any Markov process-based training that has a Gibbs-style stationary distribution. The proof is surprisingly simple, once we observe that the marginal distribution divergence from initialization remains bounded, as implied by a generalized second law of thermodynamics.
@@ 1 Introduction
It is still unclear why overparameterized models, such as neural networks (NN), typically generalize well in practice — even though there exist parameters that fit the training set but generalize poorly [20, 54]. Many works aimed to derive generalization bounds that explain this behavior. However, the generalization mystery in such overparameterized nonconvex models is still far from being “solved”.
In this work, we derive novel generalization results for two common types of stochastic processes:
  • In Section 2 we develop, for the first time, a generalization bound applicable to any data-dependent Markov process with a Gibbs-type stationary distribution (i.e. whose finite density exists and is nonzero w.r.t. some data-independent base measure). This bound (Theorem 2.7) is applicable at any time, regardless of time or the trajectory of the process.
  • In Section 3 we focus on the common special case of Continuous-time Langevin Dynamics (CLD). Our bound (Corollary 3.1) can provide a generalization result, which is, for the first time, independent of dimension, number of iterations, or the trained weights. This is in contrast to existing bounds for CLD (see Section 4.1 and Table 1), where these dependencies cause the bounds to become less useful in practical scenarios such as NN training (e.g. be vacuous).
The proofs build on a key technical result (Corollary 2.5), showing how for the marginal state distribution, its divergence (either KL or the Rényi infinity divergence) from initialization is bounded due to its monotonicity, i.e. a generalized second law of thermodynamics (see 10, 38).
@@ 2 Generalization Bounds for General Markov Process
In this section, we focus on the main ideas and results. For the detailed proofs of all the results in this section, see Appendix B. Specifically, we consider general data-dependent Markov processes over predictors and obtain a bound on their generalization gap. Importantly, although the bound only depends on the initialization distribution and a stationary distribution, it will apply to predictors at any time $t\geq 0$ along the Markov process. Our main goal is to apply these bounds to stochastic training methods, such as Langevin dynamics, where the iterates form a data-dependent Markov process. But to emphasize the broad generality of the results, in this section we consider a generic stochastic optimization framework and general data-dependent Markov processes.
We obtain generalization guarantees by bounding the KL-divergence (or, for high probability bounds, the Rényi infinity divergence) between the data-dependent marginal state distribution $p_{t}$ of the predictors at time $t$, and some data-independent base measure $\nu$ (the PAC-Bayes “prior”). The crux of the analysis is therefore bounding the divergence between $p_{t}$ and $\nu$, based only on assumptions on the initial distribution $p_{0}$ (specifically, the divergence between $p_{0}$ and $\nu$) and a stationary distribution $p_{\infty}$ (specifically, requiring that $p_{\infty}$ can be expressed as a Gibbs distribution with bounded potential or expected potential) — we do this in Section 2.1. Then, in Section 2.2 we plug these bounds on the divergence between $p_{t}$ and $\nu$ into standard PAC-Bayes bounds to obtain the desired generalization guarantees.
@@ 2.1 Bounding the Divergence of a Markov Process
Divergences and Gibbs distributions. We recall the definitions of our two divergences, and also relate them to the Gibbs distribution. It will also be convenient for us to introduce “relative” versions of divergences. In what follows, for measures $p,q$ we denote by $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$ the Radon-Nikodym derivative (i.e. the density of $p$ w.r.t. $q$) when it exists (this happens when $p\ll q$, i.e. $p$ is absolutely continuous w.r.t. $q$), or $\infty$ otherwise.
Definition 2.1 (Divergences).
@@
    The $\mu$-weighted Kullback-Leibler (KL) divergence (a.k.a. relative cross-entropy) is $\mathrm{KL}_{\mu}\left(p\,\middle\|\,q\right)=\int\mathop{}\!\mathrm{d}\mu\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}$, and the KL-divergence is then $\mathrm{KL}\left(p\,\middle\|\,q\right)=\mathrm{KL}_{p}\left(p\,\middle\|\,q\right)$.
@@
    The Rényi infinity divergence is $D_{\infty}^{\mu}\left(p\,\|\,q\right)=\operatorname{ess\,sup}_{\mu}\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}$, with $D_{\infty}\left(p\,\|\,q\right)=D_{\infty}^{p}\left(p\,\|\,q\right)$.
@@
That is, the potential of a Gibbs distribution $p$ allows us to bound the divergence in both directions between $p$ and the base measure $q$. A generalized converse of Claim 2.3 also holds, and we have that bounding on the symmetrized divergences (but not just on one direction!) is also sufficient for $p$ being Gibbs with a bounded potential. [^2] More formally: $\mathrm{KL}\left(p\,\middle\|\,q\right)+\mathrm{KL}\left(q\,\middle\|\,p\right)\leq\beta$ iff there exists a potential $\Psi$ such that $p$ is Gibbs w.r.t. $q$ with potential $\Psi$ and $\mathbb{E}_{q}\Psi-\mathbb{E}_{p}\Psi\leq\beta$, and similarly $D_{\infty}\left(p\,\|\,q\right)+D_{\infty}\left(q\,\|\,p\right)\leq\beta$ iff there exists a potential $0\leq\Psi\leq\beta$ such that $p$ is Gibbs w.r.t. $q$ with potential $\Psi$. See Claim B.9 for a proof.
@@
Let $p_{t}$ be the state distribution of a time-invariant Markov process, and $p_{\infty}$ a stationary distribution for the transitions of the Markov process (the process need not be ergodic, and $p_{t}$ need not converge to $p_{\infty}$). Then for any $t\geq 0$
@@
When the stationary distribution is uniform (thus having maximal entropy), the KL-form of Claim 2.4 recovers the familiar second law of thermodynamics, i.e. that the entropy is monotonically non-decreasing. The more general form, as in Claim 2.4, is a direct consequence of the data processing inequality, as pointed out by Cover [10] (see also [11, 38] and the generalization to Rényi divergences in [50, Theorem 9 and Example 2] —for completeness we provide a proof in Section A.2).
In our case, the stationary distribution $p_{\infty}$ will not be uniform, but rather will be very data-dependent (we are interested mostly in processes that aim to optimize some data-dependent quantity, such as Langevin dynamics). Nevertheless, we do want to use Claim 2.4 to control the entropy of $p_{t}$ relative to some benign data-independent base distribution $\nu$ (which we can informally think of as “uniform”). To do so, we can use the chain rule and plug in Claim 2.4 to obtain that for any distribution $\nu$ and at any time $t$ we have (see Lemma B.4 in Appendix B for the full derivation):
@@
Bounding the last two terms in (1) and (2) using Claim 2.3 we obtain the main result of this subsection:
@@
The important feature of Corollary 2.5 is that it bounds the divergence at any time $t$, in terms of a right-hand side that depends only on the initial distribution $p_{0}$ and a stationary distribution $p_{\infty}$. Interpreting the divergence $D\left(p_{t}\,\|\,\nu\right)$ as a measure of concentration, the Corollary ensures that at no point during its run, and regardless of mixing, does the Markov process concentrate too much, and it always maintains high entropy (relative to the base measure $\nu$).
@@
In order to bound the divergence $D\left(p_{t}\,\|\,\nu\right)$ at finite time $t$, it is not enough to rely only on the divergences $D\left(p_{0}\,\|\,\nu\right)$ and $D\left(p_{\infty}\,\|\,\nu\right)$ from the initial and stationary distributions, and it is necessary to rely also on the reverse divergence $D\left(\nu\,\|\,p_{\infty}\right)$ — see Appendix C.
@@ 2.2 From Divergences to Generalization
Corollary 2.5 can be directly used to obtain PAC-Bayes type generalization guarantees. Specifically, we consider a generic stochastic optimization setting specified by a bounded instantaneous objective $f:\mathcal{H}\times\mathcal{Z}\rightarrow[0,1]$ over a class $\mathcal{H}$, which we will refer to as the “predictor” class, and instance domain $\mathcal{Z}$. For example, in supervised learning $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and $f(h,(x,y))=\mathbb{I}\left\{h(x)\neq y\right\}$ measures the error of predicting $h(x)$ when the correct label is $y$. For a source distribution $D$ over $\mathcal{Z}$ and data $S\sim D^{N}$ of size $N$ we would like to relate the population and empirical objectives
@@
We present two types of generalization grantees: grantees that hold in expectation over a draw from the Markov process ((6) below) and grantees that hold with high probability over a single draw from the Markov process (as in (7), e.g. a single run of CLD). In both cases, the guarantees hold with high probability over the training set.
@@
Consider any distribution $D$ over $\mathcal{Z}$, function $f:\mathcal{H}\times\mathcal{Z}\to[0,1]$, and sample size $N>1$, any distribution $\nu$ over $\mathcal{H}$, and any discrete or continuous time process $\{h_{t}\in\mathcal{H}\}_{t\geq 0}$ (i.e. $t\in\mathbb{Z}_{+}$ or $t\in\mathbb{R}_{+}$) that is time-invariant Markov conditioned on $S$. Denote $p_{0}(\cdot;S)$ the initial distribution of the Markov process (that may depend on $S$). Let $p_{\infty}(\cdot;S)$ be any stationary distribution of the process conditioned on $S$, and $\Psi_{S}(h)\geq 0$ a non-negative potential function that can depend arbitrarily on $S$, such that $p_{\infty}(\cdot;S)$ is Gibbs w.r.t. $\nu$ with potential $\Psi_{S}$. Then: 1. with probability $1-\delta$ over $S\sim D^{N}$, $\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]$ $\leq\sqrt{\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{1}}{{\delta}}}{N}}\,,$ (6) 2. with probability $1-\delta$ over $S\sim D^{N}$ and over $h_{t}$: $E_{D}(h_{t})-E_{S}(h_{t})$ $\leq\sqrt{\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{h\sim p_{0}(\cdot;S)}\Psi_{S}(h)+\ln\nicefrac{{1}}{{\delta}}}{N}}\,.$ (7)
@@
The Theorem follows immediately by plugging the divergence bounds of Corollary 2.5 into standard PAC-Bayes guarantees, reproduced from Alquier et al. [1] as Theorem B.1 (in expectation, for part 1) and Theorem B.2 (single-sample, for part 2) in Appendix B. Note that we can just as easily plug Corollary 2.5 into tighter, though less explicit, PAC-Bayes bounds (see Appendix B). ∎
@@
A simplified variant of Theorem 2.7 can be stated when the initial distribution $p_{0}$ is data-independent and always equal to $\nu$. In this case the divergence between $p_{0}$ and $\nu$ vanishes, and (6) and (7) become
$\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]\leq\sqrt{\frac{\mathbb{E}_{p_{0}}\Psi_{S}+\ln\nicefrac{{1}}{{\delta}}}{N}},\quad E_{D}(h_{t})-E_{S}(h_{t})\leq\sqrt{\frac{\operatorname{ess\,sup}_{p_{0}}\Psi_{S}+\ln\nicefrac{{1}}{{\delta}}}{N}}\,.$ (8)
@@
In Theorem 2.7, the Markov process need not be ergodic, and need not converge to $p_{\infty}$, or converge at all. If there are multiple stationary distributions, the theorem holds for all of them, and so we can take $p_{\infty}$ to be any stationary distribution we want. And in any case, there is no mixing requirement, and the theorem holds at any time $t$.
@@
Our data-dependent Markov process of interest, and in particular CLD and SGD, might aim to minimize $E_{S}(h_{t})$, and the potential $\Psi$ might also be related to it (as in, e.g., CLD). This is allowed, but is in no way required in Theorem 2.7. Even for CLD, these might be related but not the same, as we might be minimizing a surrogate loss, such as a logistic loss, but are interested in bounding the generalization gap for a zero-one error. In stating Theorem 2.7 we intentionally refer to an arbitrary stochastic optimization problem and an arbitrary data-dependent Markov process, that are allowed to be related or dependent in arbitrary ways.
@@
In Appendix C we show that in order to ensure generalization at every intermediate $t$, it is not sufficient to only bound $\mathrm{KL}\left(p_{\infty}\,\middle\|\,\nu\right)$ or $D_{\infty}\left(p_{\infty}\,\|\,\nu\right)$, and we do need the stronger symmetric bound ensured by the Gibbs potential and Claim 2.3; and that it is also necessary to relate both $p_{0}$ and $p_{\infty}$ to the same data independent distribution $\nu$, as relating them to different data-independent distributions ensures generalization at the beginning and at the end, but not the middle of training.
@@
In Theorem 2.7 we plugged Corollary 2.5 into a simple PAC-Bayes bound, providing a uniform bound on the generalization gap that allows for easy interpretation and comparison with other results. But once we have the divergence bounds of Corollary 2.5, we can just as easily plug them into tighter PAC-Bayes bounds—see Appendix B. For example, when $\mathbb{E}\left[E_{S}\left(h_{t}\right)\mid S\right]\approx 0$, we obtain there a $O\left(1/N\right)$ rate, which is faster than the $O\left(1/\sqrt{N}\right)$ rate of Theorem 2.7.

### v2

@@ Abstract
We analyze the generalization gap (gap between the training and test errors) when training a potentially over-parametrized model using a Markovian stochastic training algorithm, initialized from some distribution ${\boldsymbol{\theta}}_{0}\sim p_{0}$. We focus on Langevin dynamics with a positive temperature $\beta^{-1}$, i.e. gradient descent on a training loss $L$ with infinitesimal step size, perturbed with $\beta^{-1}$-variances Gaussian noise, and lightly regularized or bounded. There, we bound the generalization gap, at any time during training, by $\sqrt{(\beta\mathbb{E}L({\boldsymbol{\theta}}_{0})+\ln(1/\delta))/N}$ with probability $1-\delta$ over the dataset, where $N$ is the sample size, and $\mathbb{E}L({\boldsymbol{\theta}}_{0})=O(1)$ with standard initialization scaling. In contrast to previous guarantees, we have no dependence on either training time or reliance on mixing, nor a dependence on dimensionality, gradient norms, or any other properties of the loss or model. This guarantee follows from a general analysis of any Markov process-based training that has a Gibbs-style stationary distribution. The proof is surprisingly simple, once we observe that the marginal distribution divergence from initialization remains bounded, as implied by a generalized second law of thermodynamics.
@@ 1 Introduction
One main goal of contemporary machine learning theory is to predict a model’s behavior before training occurs. A commonly desired metric is the generalization of overparameterized models, such as neural networks (NN). For these models, such a predictive theory of generalization is still lacking, despite great empirical success [71, 23]. In particular, a significant line of work aimed to explain the role of optimization in generalization (e.g. [23, 64, 40, 66]), and specifically the effect of stochasticity (e.g. [59, 49, 10, 8]).
Data-dependent Markov processes are a common optimization approach. These include stochastic gradient descent (SGD), as well as other stochastic gradient methods either studied theoretically [30, 59], or used in practice such as SGD with momentum [52], ADAM [34], and many more. Of particular interest are continuous Langevin dynamics (CLD) and discrete analogues of it, which have been studied extensively as models for SGD (see Section˜4.1).
In Section˜2 we develop, for the first time, a generalization bound applicable to any data-dependent Markov process with a Gibbs-type stationary distribution (i.e. whose finite density exists and is nonzero w.r.t. some data-independent base measure). An important feature of our analysis is that it is entirely independent of the training time $t$, both in that we do not rely on training for only a small number of steps, nor that we rely on mixing — the guarantees are valid at any time, with no dependence at all on $t$. Furthermore, it is also completely trajectory independent.
In Section˜3 we apply these general results to the particular case where training is done with CLD with loss $L$ and inverse temperature $\beta$, deriving a particularly simple generalization bound for CLD, which we compare to previous generalization bounds for CLD in Section˜4, as well as discussing other related work. Finally, we address limitations and future work in Section˜5.
To prove these results, we first show in Section˜2 how, for the marginal distribution at time $t$, $p_{t}$, its divergence (either KL or the Rényi infinity divergence) from initialization is bounded due to its monotonicity, i.e. a generalized second law of thermodynamics [11, 46]. This surprisingly simple derivation111e.g. to bound the KL divergence of a Markov process having a stationary distribution with potential $\Psi\in\left[0,\infty\right)$, i.e. $\mathop{}\!\mathrm{d}p_{\infty}/\mathop{}\!\mathrm{d}p_{0}\propto e^{-\Psi}$ (e.g., $\Psi=\beta L$ for CLD), the second law implies the first inequality below $\mathrm{KL}\!\left(p_{t}||p_{0}\right)$ $=\!\int\!p_{t}\ln\frac{p_{t}}{p_{0}}\!=\!\int\!p_{t}\ln\frac{p_{t}}{p_{\infty}}\!+\!\int\!p_{t}\ln\frac{p_{\infty}}{p_{0}}\!\leq\!\int\!p_{0}\ln\frac{p_{0}}{p_{\infty}}\!+\!\int\!p_{t}\ln\frac{p_{\infty}}{p_{0}}\!=E_{p_{0}}\Psi-E_{p_{t}}\Psi\leq E_{p_{0}}\Psi.$ leads to our key technical result (Corollary˜2.5). Standard PAC-Bayes generalization bounds [43] then yield our generalization bounds (Theorems˜2.7 and 3.1).
@@ 2 Generalization Bounds for General Markov Process
In this Section, we consider general data-dependent Markov processes over predictors and obtain a bound on their generalization gap. Importantly, although the bound only depends on the initialization distribution and a stationary distribution, it will apply to predictors at any time $t\geq 0$ along the Markov process. Our main goal is to apply these bounds to stochastic training methods, such as Langevin dynamics, where the iterates form a data-dependent Markov process. But to emphasize the broad generality of the results, in this section we consider a generic stochastic optimization framework and general data-dependent Markov processes.
We obtain generalization guarantees by bounding the KL-divergence (or, for high probability bounds, the Rényi infinity divergence, see Definition˜2.1) between the data-dependent marginal distribution $p_{t}$ of the predictors at time $t$, and some data-independent base measure $\nu$ (the PAC-Bayes “prior”). The crux of the analysis is therefore bounding the divergence between $p_{t}$ and $\nu$, based only on assumptions on the initial distribution $p_{0}$ (specifically, the divergence between $p_{0}$ and $\nu$) and a stationary distribution $p_{\infty}$ (specifically, requiring that $p_{\infty}$ can be expressed as a Gibbs distribution with bounded potential or expected potential, see Definition˜2.2) — we do this in Section˜2.1. Then, in Section˜2.2 we plug these bounds on the divergence between $p_{t}$ and $\nu$ into standard PAC-Bayes bounds to obtain the desired generalization guarantees.
Detailed proofs of all the results in this section can be found in Appendix˜B.
@@ 2.1 Bounding the Divergence of a Markov Process
Divergences and Gibbs distributions. We recall the definitions of our two divergences, and also relate them to the Gibbs distribution. It will also be convenient for us to introduce “relative” versions of divergences.
Definition 2.1 (Divergences  [^3] The term “divergence” is a slight abuse of notation, as the following definitions are not strictly non-negative, without specifying $\mu$.).
@@
    The $\mu$-weighted Kullback-Leibler (KL) divergence (a.k.a. relative cross-entropy) is [^4] For two measures $p$ and $q$, $\mathop{}\!\mathrm{d}p/\mathop{}\!\mathrm{d}q$ is the Radon-Nikodym derivative (i.e. the density of $p$ w.r.t. $q$) when it exists (i.e. when $p\ll q$, i.e. $p$ is absolutely continuous w.r.t. $q$), or $\infty$ otherwise. $\mathrm{KL}_{\mu}\left(p\,\middle\|\,q\right)=\int\mathop{}\!\mathrm{d}\mu\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}$, and the KL-divergence is then $\mathrm{KL}\left(p\,\middle\|\,q\right)=\mathrm{KL}_{p}\left(p\,\middle\|\,q\right)$.
@@
    The Rényi infinity divergence is [^5] The essential supremum of a function $f$ w.r.t. a measure $\mu$ is $\operatorname{ess\,sup}_{\mu}f=\inf\left\{b\in\mathbb{R}\,\mid\,\mu\left(f>b\right)=0\right\}$, i.e. the smallest (infimum) number that bounds $f$ from above almost everywhere. The essential infimum is defined similarly. $D_{\infty}^{\mu}\left(p\,\|\,q\right)=\operatorname{ess\,sup}_{\mu}\ln\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}q}$, with $D_{\infty}\left(p\,\|\,q\right)=D_{\infty}^{p}\left(p\,\|\,q\right)$.
@@
That is, the potential of a Gibbs distribution $p$ allows us to bound the divergence in both directions between $p$ and the base measure $q$. A generalized converse of Claim˜2.3 also holds, and we have that bounding on the symmetrized divergences (but not just on one direction!) is also sufficient for $p$ being Gibbs with a bounded potential. [^6] More formally: $\mathrm{KL}\left(p\,\middle\|\,q\right)+\mathrm{KL}\left(q\,\middle\|\,p\right)\leq\beta$ iff there exists a potential $\Psi$ such that $p$ is Gibbs w.r.t. $q$ with potential $\Psi$ and $\mathbb{E}_{q}\Psi-\mathbb{E}_{p}\Psi\leq\beta$, and similarly $D_{\infty}\left(p\,\|\,q\right)+D_{\infty}\left(q\,\|\,p\right)\leq\beta$ iff there exists a potential $0\leq\Psi\leq\beta$ such that $p$ is Gibbs w.r.t. $q$ with potential $\Psi$. See Claim B.8 for a proof.
@@
Let $p_{t}$ be the marginal distribution of a time-invariant Markov process, and $p_{\infty}$ a stationary distribution for the transitions of the Markov process (the process need not be ergodic, and $p_{t}$ need not converge to $p_{\infty}$). Then for any $t\geq 0$
@@
When the stationary distribution is uniform (thus having maximal entropy), the KL-form of Claim˜2.4 recovers the familiar second law of thermodynamics, i.e. that the entropy is monotonically non-decreasing. The more general form, as in Claim˜2.4, is a direct consequence of the data processing inequality, as pointed out by Theorem 4 of Cover [11] (see also [12, 46] and the generalization to Rényi divergences in [65, Theorem 9 and Example 2] —for completeness we provide a proof in Section˜A.2).
In our case, the stationary distribution $p_{\infty}$ will not be uniform, but rather will be very data-dependent (we are interested mostly in processes that aim to optimize some data-dependent quantity, such as Langevin dynamics). Nevertheless, we do want to use Claim˜2.4 to control the entropy of $p_{t}$ relative to some benign data-independent base distribution $\nu$ (which we can informally think of as “uniform”). To do so, we can use the chain rule and plug in Claim˜2.4 to obtain that for any distribution $\nu$ and at any time $t$ we have (see Lemma˜B.1 in Appendix˜B for the full derivation):
@@
Bounding the last two terms in (1) and (2) using Claim˜2.3 we obtain the main result of this subsection:
@@
The important feature of Corollary˜2.5 is that it bounds the divergence at any time $t$, in terms of a right-hand side that depends only on the initial distribution $p_{0}$ and a stationary distribution $p_{\infty}$. Interpreting the divergence $D\left(p_{t}\,\|\,\nu\right)$ as a measure of concentration, the Corollary ensures that at no point during its run, and regardless of mixing, does the Markov process concentrate too much, and it always maintains high entropy (relative to the base measure $\nu$).
@@
In order to bound the divergence $D\left(p_{t}\,\|\,\nu\right)$ at finite time $t$, it is not enough to rely only on the divergences $D\left(p_{0}\,\|\,\nu\right)$ and $D\left(p_{\infty}\,\|\,\nu\right)$ from the initial and stationary distributions, and it is necessary to rely also on the reverse divergence $D\left(\nu\,\|\,p_{\infty}\right)$ — see Appendix˜C.
@@ 2.2 From Divergences to Generalization
Corollary˜2.5 can be directly used to obtain PAC-Bayes type generalization guarantees. Specifically, we consider a generic stochastic optimization setting specified by a bounded instantaneous objective $f:\mathcal{H}\times\mathcal{Z}\rightarrow[0,1]$ over a class $\mathcal{H}$, which we will refer to as the “predictor” class, and instance domain $\mathcal{Z}$. For example, in supervised learning $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and $f(h,(x,y))=\mathbb{I}\left\{h(x)\neq y\right\}$ measures the error of predicting $h(x)$ when the correct label is $y$. For a source distribution $D$ over $\mathcal{Z}$ and data $S\sim D^{N}$ of size $N$ we would like to relate the population and empirical objectives
@@
We present two types of generalization guarantees: guarantees that hold in expectation over a draw from the Markov process ((6) below) and guarantees that hold with high probability over a single draw from the Markov process (as in (7), e.g. a single run of CLD). In both cases, the guarantees hold with high probability over the training set.
@@
Consider any distribution $D$ over $\mathcal{Z}$, function $f:\mathcal{H}\times\mathcal{Z}\to[0,1]$, sample size $N\geq 8$, and any distribution $\nu$ over $\mathcal{H}$. Let $\{h_{t}\in\mathcal{H}\}_{t\geq 0}$ be a discrete or continuous time process (i.e. $t\in\mathbb{Z}_{+}$ or $t\in\mathbb{R}_{+}$) that is time-invariant Markov conditioned on $S$, that starts from an initial distribution $p_{0}(\cdot;S)$ (that may depend on $S$), and admits a stationary distribution conditioned on $S$, $p_{\infty}(\cdot;S)$. Let $\Psi_{S}(h)\geq 0$ be a non-negative potential function and assume that $p_{\infty}(\cdot;S)$ is Gibbs w.r.t. $\nu$ with potential $\Psi_{S}$. Then: 1. with probability $1-\delta$ over $S\sim D^{N}$, $\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]$ $\leq\sqrt{\frac{\mathrm{KL}\left(p_{0}(\cdot;S)\,\middle\|\,\nu\right)+\mathbb{E}\left[\Psi_{S}(h_{0})\middle|S\right]+\ln\nicefrac{{N}}{{\delta}}}{2N}}\,$, (6) 2. with probability $1-\delta$ over $S\sim D^{N}$ and over $h_{t}$: $E_{D}(h_{t})-E_{S}(h_{t})$ $\leq\sqrt{\frac{D_{\infty}\left(p_{0}(\cdot;S)\,\|\,\nu\right)+\operatorname{ess\,sup}_{h\sim p_{0}(\cdot;S)}\Psi_{S}(h)+\ln\nicefrac{{N}}{{\delta}}}{2N}}\,.$ (7)
@@
The Theorem follows immediately by plugging the divergence bounds of Corollary˜2.5 into standard PAC-Bayes guarantees, which we do in Appendix˜B. ∎
@@
A simplified variant of Theorem˜2.7 can be stated when the initial distribution $p_{0}$ is data-independent and always equal to $\nu$. In this case the divergence between $p_{0}$ and $\nu$ vanishes, and (6) and (7) become
$\mathbb{E}\left[E_{D}(h_{t})-E_{S}(h_{t})\middle|S\right]\leq\!\sqrt{\frac{\mathbb{E}_{p_{0}}\left[\Psi_{S}\mid S\right]+\ln\nicefrac{{N}}{{\delta}}}{2N}},\;E_{D}(h_{t})-E_{S}(h_{t})\leq\!\sqrt{\frac{\operatorname{ess\,sup}_{p_{0}}\Psi_{S}+\ln\nicefrac{{N}}{{\delta}}}{2N}}\,.$ (8)
@@
In Theorem˜2.7, the Markov process need not be ergodic, and need not converge to $p_{\infty}$, or converge at all. If there are multiple stationary distributions, the theorem holds for all of them, and so we can take $p_{\infty}$ to be any stationary distribution we want. And in any case, there is no mixing requirement, and the theorem holds at any time $t$.
@@
Our data-dependent Markov process of interest, and in particular CLD and SGD, might aim to minimize $E_{S}(h_{t})$, and the potential $\Psi$ might also be related to it (as in, e.g., CLD). This is allowed, but is in no way required in Theorem˜2.7. Even for CLD, these might be related but not the same, as we might be minimizing a surrogate loss, such as a logistic loss, but are interested in bounding the generalization gap for a zero-one error. In stating Theorem˜2.7 we intentionally refer to an arbitrary stochastic optimization problem and an arbitrary data-dependent Markov process, that are allowed to be related or dependent in arbitrary ways.
@@
In Appendix˜C we show that in order to ensure generalization at every intermediate $t$, it is not sufficient to only bound $\mathrm{KL}\left(p_{\infty}\,\middle\|\,\nu\right)$ or $D_{\infty}\left(p_{\infty}\,\|\,\nu\right)$, and we do need the stronger symmetric bound ensured by the Gibbs potential and Claim˜2.3; and that it is also necessary to relate both $p_{0}$ and $p_{\infty}$ to the same data independent distribution $\nu$, as relating them to different data-independent distributions ensures generalization at the beginning and at the end, but not the middle of training.
@@
In Theorem˜2.7 we plugged Corollary˜2.5 into a simplified PAC-Bayes bound that allows for easy interpretation and comparison with other results. But once we have the divergence bounds of Corollary˜2.5, we can just as easily plug them into tighter PAC-Bayes bounds — see Appendix˜B. For example, when $E_{S}\left(h_{t}\right)\approx 0$, these yield a rate of $O\left(1/N\right)$.